OpenStudy (anonymous):

sin(-pi)+cos(5pi) Find the exact value

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sin(-\pi)~~\Rightarrow~~-\sin(\pi)}\) Do yo know what \(\large\color{black}{ \displaystyle \sin(\pi)}\) is ?

OpenStudy (agl202):

adding (or subtracting) 2pi inside these trig functions takes you once around the unit circle and back to the same point. sin (-pi) = sin(pi) = 0 cos (5pi) = cos (3pi) = cos (pi) =-1

OpenStudy (solomonzelman):

sin(-A)=-sin(A) cos(-A)=cos(A)

OpenStudy (solomonzelman):

it is -sin(pi) which is technically the same thing, but just not to apply the rules for sines of other values

OpenStudy (solomonzelman):

And then \(\normalsize\color{black}{ \displaystyle \cos(5\pi)~~\Rightarrow~~\cos(4\pi+\pi)~~\Rightarrow~~\cos(4\pi)\cos(\pi)-\sin(4\pi)\sin(\pi) }\) sin(pi) and cos(pi) should be the values you are familiar with.

OpenStudy (anonymous):

I think I get it

OpenStudy (solomonzelman):

that is cooL:)